Strassel Investors buys real estate, develops it, and resells it for a profit. A new property is available, and Bud Strassel, the president and owner of Strassel Investors, believes if he purchases and develops this property it can then be sold for $160,000. The current property owner has asked for bids and stated that the property will be sold for the highest bid in excess of $100,000. Two competitors will be submitting bids for the property. Strassel does not know what the competitors will bid, but he assumes for planning purposes that the amount bid by each competitor will be uniformly distributed between $100,000 and $150,000.

  1. Develop a worksheet that can be used to simulate the bids made by the two competitors. Strassel is considering a bid of $130,000 for the property. Using a simulation of 1000 trials, what is the estimate of the probability Strassel will be able to obtain the property using a bid of $130,000? Round your answer to 1 decimal place. Enter your answer as a percent.

    _________%
  2. How much does Strassel need to bid to be assured of obtaining the property?

    $ __________

    What is the profit associated with this bid?

    $ __________
  3. Use the simulation model to compute the profit for each trial of the simulation run. With maximization of profit as Strassel’s objective, use simulation to evaluate Strassel’s bid alternatives of $130,000, $140,000, or $150,000. What is the recommended bid, and what is the expected profit?

    A bid of __________ results in the largest mean profit of $ _________ .
Answer #1

The simulation model is following:

FORMULAS:

B9 =100000+RAND()*(150000-100000) copy to B9:C1008

D9 =MAX(B9:C9) copy to D9:D1008

E9 =IF(D9<$C$3,1,0) copy to E9:E1008

C5 =AVERAGE(E9:E1008)

a) Probability = 36.7 %

b) To be assured of obtaining the property, Strassel must bid = $ 150,000

Profit associated with this bid = 160,000 - 150,000 = $ 10,000

c) The simulation model for profit maximization is following:

A new field for Profit is added in the earlier model.

Formula for profit: F9 =(160000-$C$3)*E9 copy to F9:F1008

C5 =AVERAGE(F9:F1008)

Using trial and error, note the average profit for bid value of 130000, 140000 and 150000

From the above simulation results, it is clear that the highest profit is achieved with a bid value fo $ 140,000

ANSWER $ 140,000

Largest mean profit = $ 13,080